Fermi National Laboratory
 

Accelerator Update for June 13, 2007 - June 15, 2007

 

Wednesday June 13

The day shift began with Operations monitoring Tevatron (TeV) store 5489 (40.92E30) and the Recycler stash (296.3E10), with the Antiproton Source (Pbar) stacking (38.7E10), with the Linac, Booster, and Main Injector (MI) operating normally, with NuMI and MiniBooNE taking beam, and with D-Zero (D0) and the Collider Detector at Fermilab (CDF) using the colliding beams.

At 11:45 AM, Operations reported that a MI kicker (I:KPS1A) had tripped off three times during this shift. It tripped off two more time over the next 17 minutes. It reset okay.

A MI expert began a slipstacking study at 11:57 AM.

At 12:27 PM, the Linac Buncher tripped off. It reset.

At 12:56 PM, two MI power supplies (North and Center Anode) tripped off.

At 1:07 PM, the MI power supply interlock loop (PSLOOP) tripped off on a “Volts to Ground” indication. It reset as well as the power supplies.

A Pbar expert began a one-hour reverse proton study at 2:02 PM.

At 2:35 PM, KPS1A tripped off. It reset.

Operations terminated store 5489 at 2:36 PM.

Operations had CDF and D0 prepared for an access by 2:48 PM.

At 3:36 PM, Operations reported that the MiniBooNE beam permit tripped off due to bad beam positions. It reset.

Operations reported at 3:56 PM that the Recycler had stashed antiprotons three times during the day shift.

D0 and CDF were both out of access by 5:09 PM.

Operations put the TeV through a dry squeeze at 5:28 PM.

TeV experts began a study period around 6 PM.

Operations established store 5492 at 10:24 PM with an initial luminosity of 245.31E30.

Operations reported at 11:50 PM that the Recycler had stashed once during the evening shift.

Thursday June 14

The midnight shift began with Operations monitoring store 5492 (183.31E30) and the stash (25.9E10), with Pbar stacking (128.9E10), and with NuMI and MiniBooNE taking beam.

At 12:11 AM, KPS1A tripped off. It reset.

Operations took the master beam switch (beam shut off) at 6 AM in preparation of the shutdown for MI LCW leak repairs.

Operations turned off Booster, MI, and the Pbar Pre-Vault area at 6:05 AM.

Kicker experts began work on the MI-10 kickers at 6:16 AM.

RF experts turned off the Booster RF for work at 6:16 AM.

Experts began work on the NuMI Lambertson (I:LAM60) at 6:20 AM.

Linac experts turned of KRF3 at 7:12 AM to replace its cavity LCW flow meter.

At 7:48 AM, RF experts reported that they found two MI RF cavity LCW leaks; the MIRF12 leak would be easy to repair. The MIRF13 leak was found to be internal and would require most of the day to replace equipment.

Operations reported at 8 AM that Recycler experts conducted studies through out most of the midnight shift. Accelerator status: store 5492 was at 91.81E30 and the stack was at 205.9E10. The MI/Recycler tunnels were in access.

Linac experts began work on KLRF2 at 8:24 AM.

Kicker experts reported at 11:41 AM, that they had completed their work on the MI-10 kickers.

At 12:48 PM, LRF4 tripped off for the first of four times in 15 minutes. It reset.

Linac experts began changing LRF4 switch tubes at 1:45 PM.

Operations brought MiniBooNE on line at 2:32 PM.

Operations brought Booster on line at 2:48 PM.

At 2:52 PM, store 5492 quenched at sector E1. Later investigations revealed that the current of a quadrupole power supply (BQ9) wandered away from its nominal state.

Operations prepared D0 and for an access at 3:52 PM.

Power supply experts checked out BQ9, but didn’t find any problems. They replaced the firing circuit module and swapped out the control cards at 4:43 PM.

MI RF experts reported around 6:45 PM that they had completed the leak repairs on MIRF13.

Linac experts reported at 5:39 PM, that they had completed their work on LRF4.

Operations established beam to NuMI and MiniBooNE at 7:34 PM.

At 7:48 PM, FESS techs reported that they had finished cleaning the TeV cryo SSB strainers, but were still trying to unclog a discharge line elbow.

At 8:09 PM, Operations reported that a Pbar quadrupole power supply (I:Q703) wasn’t ramping. A Pbar expert suggested that operators swap out a program card, but that didn’t help. Operators called in a power supply expert.

At 8:36 PM, KPS1A tripped off. It reset.

At 10:10 PM, KRF6 tripped off and wouldn’t reset. Operators called in an expert.

A power supply expert reported at 10:20 PM that he replaced the original DAC card for I:Q703 and the supply resumed normal operations.

Pbar resumed stacking at 10:50 PM, but then emittances went bad and half the stack was lost. Operators investigated and then called in an expert.

Operations began ramping the TeV at 11:07 PM, but the sector A2 cryo system started to go unstable. Operators halted the ramp and contacted experts.

At 11:37 PM, Operations reported that KRF6 had tripped off for the first of four times over the next hour. Experts investigated and adjusted the switch tube level and swapped the CS SWBalance card.

Operations turned on the Pelletron at 11:53 PM to verify it was operating normally. It was then turned off.

At midnight, Operations reported the status of work and of the machines during the evening shift:

Proton source

  • Booster ready for beam at start of shift
  • Linac KRF7 repaired (fuse in relay contactor replaced) by 5 PM
  • Linac LRF4 switch tubes replaced by 5:30 PM
  • Beam to Booster dump established by 5:30 PM

Main Injector

  • MIRF water leak repaired by 6:45 PM
  • Beam to NuMI & MiniBooNE by 7:30 PM
  • Protons to RR by 7:45 PM
  • I:Q703 not ramping 8 – 10:15 PM. Original DAC card put back

Pbar

  • Pbar idle holding 200mA stash until 10:45 PM
  • As beam to Pbar was established, a transverse instability knocked out 125 mA, leaving 75mA in the stack
  • Experts diagnosed problems into the owl shift

Tevatron

  • C:BQ9 (cause of day shift quench) firing circuit swap done by 4:30 PM
  • SSB cryo water system strainers cleaned 4 – 10:30 PM
  • A0 cryo water system pump repaired 5 - 5:15 PM
  • Current on TeV bus by 8:30 PM
  • D0 controlled access 4:30 - 6:30 PM
  • HEP ramping by 11 PM
  • A2 cryo system problems held off ramping and dry squeeze thru end of shift.

Friday June 15

The midnight shift began with Operations monitoring the stash (1.2E10), with Pbar stacking (75.3E10), and with NuMI and MiniBooNE taking beam.

The A2 cryo system went stable at 12:06 AM and the Cryo Coordinator gave permission for Operations to ramp the TeV.

Operations put the TeV through a wet squeeze at 1:39 AM.

TeV experts began a study period at 2:12 AM.

Operations put the TeV through a dry squeeze 5:33 AM.

Operations reported at 7:57 AM that the Recycler had stashed antiprotons three times during the midnight shift.

The Plans for Friday and the Weekend

The plans for today and the weekend are to stack, stash, and store.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 06/15/2007    email Fermilab